OpenStudy (anonymous):

f(x) = 0.05x g(x) = (0.05)x Which statement best describes the graph of f(x) and g(x)? The graph of g(x) will eventually exceed the graph of f(x). The graph of f(x) will eventually exceed the graph of g(x). The graphs will both have their y-intercept equal to 1. The graphs will both have their y-intercept equal to 0.05

OpenStudy (anonymous):

B?

OpenStudy (freckles):

is f and g really the same function? or did you mean to write something different?

OpenStudy (anonymous):

no their two different functions

OpenStudy (freckles):

you do know that (5)x is the same as saying 5 times x which can also be written as 5x you have the same function written twice here

OpenStudy (anonymous):

f(x) = 0.05x g(x) = (0.05)^x

OpenStudy (anonymous):

oops sorry

OpenStudy (freckles):

hint \[g(x)=(0.05)^x=(\frac{1}{20})^x=20^{-x}\]

OpenStudy (freckles):

if you still don't know what to do with that graph both functions see what happens as x gets larger and larger one of the functions should be way greater than the other for really large values of x

OpenStudy (anonymous):

does f(x) become greater than g x?

OpenStudy (freckles):

yep yep in fact there is a rough graph